If ADAM is half as old as BRETT and BRETT is half as old again as CHLOE and the sum of their three ages is 105 years, find their ages
Answer by Maths68(1474)

Let
Present age of Chole = x
Present age of of Brett = x/2
Present age of Adam = x/2/2 = x/4
x+x/2+x/4=105
Multiply by LCM (4) both sides of the equation
4x+2x+x=105*4
7x=420
7x/7=420/7
x=60
Present age of Chole = x = 60 years old
Present age of of Brett = x/2 = 60/2 = 30 years old
Present age of Adam = x/2/2 = x/4 = 60/4 = 15 years old